Find values for m and b in the following system so that the solution to the system
is (-3,4).
5x+7y=b
mx+y=22
Please help!!
Thanks so much!!!

Answer by funmath(2933)

Find values for m and b in the following system so that the solution to the system
is (-3,4).
5x+7y=b
mx+y=22
:
Let x=-3 and y=4 in the first equation.
5(-3)+7(4)=b
-15+28=b
13=b
:
Let x=-3 and y=4 in the second equation.
m(-3)+4=22
-3m+4=22
-3m+4-4=22-4
-3m=18
-3m/-3=18/-3
m=-6
:
Happy Calculating!!!!